# On Mixed-Precision Iterative Methods and Analysis for Nearly Completely Decomposable Markov Processes

### v1

@@ 4.1 Error and Convergence Analysis of Conservatively Modified KMS Algorithm
Let $\mathbf{s}=\mathbf{s}\mathbf{R}$ and $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ with $\|\mathbf{s}\|_{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$, and suppose $\bm{\pi}^{(t-1)}$ is an $O(\varphi)$ approximation to $\bm{\pi}$, $t\in\mathbb{N}$. Then, we have
$\mathbf{s}^{(t-1)}-\mathbf{s}=O(\varphi),\qquad t\in\mathbb{N},$
@@
Recall $\mathbf{v}_{i}$ to be the left eigenvector of $\mathbf{P}_{ii}$ corresponding to its dominant eigenvalue $\lambda_{i_{1}}=1-{g}_{i}\epsilon+o(\epsilon)$, $i\in[m]$. We now establish our main results that shows there is an $O(\epsilon)$ norm difference between the exact stationary distribution $\bm{\pi}$ and the corresponding eigenvectors $\mathbf{v}_{i}$ of $\mathbf{P}_{ii}$ multiplied by an $O(1)$ constant and that shows Algorithm 1 is guaranteed to converge with an approximation error of the estimate $\bm{\pi}^{(t)}$ in each iteration $t$ of the exact solution $\bm{\pi}$ decreasing by a factor of $O(\epsilon)$.
@@
Let $\mathbf{v}_{i}$ be the left eigenvector of $\mathbf{P}_{ii}$ corresponding to the exact invariant probability vector $\bm{\pi}_{i}$, $i\in[m]$. Then, there exists a constant ${b}_{i}=O(1)$ such that
$\|\bm{\pi}_{i}-{b}_{i}\mathbf{v}_{i}\|_{1}=O(\epsilon).$ (16)
Moreover, Algorithm 1 converges with an approximation error in the solution $\bm{\pi}^{(t)}$ at each iteration $t$ that decreases by a factor of $O(\epsilon)$.
The results of Lemma 4.1 and Theorem 4.2 are based on the assumption that the systems of linear equations in Step $\mathbf{5}$ and Step $\mathbf{3}$ are all solved at the exact accuracy of full precision. Based on our selection of the computational precision with respect to the condition number and the norm of the matrix of the linear system as defined in Section 3.1, we next present a lemma that shows our use of the iterative refinement method in Algorithm 1 is guaranteed to converge with the approximation error of the estimate of the exact solution in each iteration decreasing by a factor of $O(\kappa(\mathbf{A}))$, where $\mathbf{A}$ is the matrix of the linear system and $\kappa(\mathbf{A})$ denotes its condition number.
@@
Consider computing the solution of the linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$ via the mixed-precision iterative refinement method under the precision selection process employed in Step $\mathbf{5}$ and Step $\mathbf{3}$ of Algorithm 1. Then, this iterative refinement method converges linearly with an approximation error that decreases by a factor of $O(\kappa(\mathbf{A}))$ in each iteration.
Convergence and related results similar to those in Lemma 4.3 hold for the alternative general classes of approximate mixed-precision computing methods of interest. We also note that the linear convergence with a decreasing factor of $O(\kappa(A))$ is typically much faster than the convergence of the outer-loop iteration of Algorithm 1 with a decreasing factor of $O(\epsilon)$.
4.2 Error and Convergence Analysis of Aggressively Modified KMS Algorithm
Our main theoretical result is that Algorithm 2 — our aggressively modified KMS algorithm — is guaranteed to converge with an approximation error that decreases with respect to a tradeoff between an increase in the number of iterations $t$ of the outer loop for fewer numbers of Richardson iteration steps $k_{t}$ in Step $\mathbf{5}$ and a decrease in the computational complexity of the Richardson iterative method for smaller $k_{t}$. To consider these theoretical aspects of our second algorithmic approach, we begin by seeking to understand the effects of stopping the mixed-precision Richardson iteration method after a certain number of steps, say $k\in\mathbb{N}$, as formally expressed in the following lemma.
Lemma 4.4.
Consider computing the solution of the linear system $(\mathbf{D}^{\top}-\mathbf{U}^{\top})\mathbf{x}=\mathbf{b}$ via the Richardson iteration method. Then, after $k$ steps, the absolute error satisfies
@@ 4.2 Error and Convergence Analysis of Aggressively Modified KMS Algorithm
Assuming that the dominant eigenvalue of the matrix $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}\mathbf{L}^{\top}\mathbf{C}$ is simple and has modulus one, the iterate $\bm{\pi}^{(t)}$ converges towards the corresponding eigenvector.
@@
Consider the update $\overline{\bm{\pi}}^{(t)\top}=((\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k_{t}]}^{-1}\mathbf{L}^{\top}\mathbf{C})\overline{\bm{\pi}}^{(t-1)\top}$ where $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}$ is applied via $k_{t}$ steps of the Richardson iteration method, and $k_{t}$ is monotonically increasing with the number of outer-loop iterations $t$. Then, the sequence $\{\overline{\bm{\pi}}^{(t)\top}\}$ converges to the same limit as the sequence $\{\bm{\pi}^{(t)\top}\}$.
@@
Suppose Algorithm 2 with exact linear system solutions provides a sequence $\{\bm{\pi}^{(t)\top}\}$ such that $\|\bm{\pi}^{(t)}-\bm{\pi}\|<\|\bm{\pi}^{(t-1)}-\bm{\pi}\|$, $t>1$. Then, the sequence $\{\overline{\bm{\pi}}^{(t)\top}\}$ produced by replacing the exact linear system solutions in Step $\mathbf{5}$ of Algorithm 2 with $k_{t}$ steps of the Richardson iteration method converges to the same limit as $\{\bm{\pi}^{(t)\top}\}$ provided that $k_{t}$ increases with the number of outer-loop iterations $t$.
Starting with an inexact application of the Richardson iteration method and increasing the number of Richardson iteration steps $k_{t}$ with the number of outer-loop iterations $t$ helps to avoid over-solving in the earlier stages of Step $\mathbf{5}$ in Algorithm 2 where the approximate stationary distribution is far from $\bm{\pi}$. For example, the tolerance up to which we solve the linear system in Step $\mathbf{5}$ follows a geometric criterion such as $\gamma^{t}$ or $\rho/(t+1)^{\theta}$ for some $\gamma,\rho\in(0,1)$ and $\theta>1$. We note here that such tradeoffs are well known in numerical linear algebra and arise often when iterative solvers are used in shift-and-invert algorithms for eigenvalue computations [12].
@@ 4.3 Mixed-Precision Computation Performance Analysis
Our general classes of algorithmic approaches exploit a combination of advances in mixed-precision computer architectures and iterative approximate computing methods. Although well studied in machine learning and related applications, we note that such advances in mixed-precision computation have received far less consideration in the context of computing the stationary distribution of Markov processes. Beyond significantly reducing computation times, our use of lower precision computation makes it possible to handle much larger systems of linear equations before hitting the memory bandwidth limitations of today’s highly-efficient advanced processors.
@@
More specifically, the performance tradeoff at the heart of our algorithmic approaches concerns, on one hand, the significant reductions in execution times afforded by mixed-precision computation at the expense of inaccuracies in the results and, on the other hand, the ability to mitigate inaccurate computations, further reduce execution times and guarantee convergence afforded by iterative approximate computing methods. Sections 4.1 and 4.2 address the degree to which our algorithmic approaches mitigate the impact of inaccurate computations, reduce execution times, and ensure convergence. The reductions in execution times provided by mixed-precision computation are also quite significant and vary from one computer processor architecture to the next, with increasing trends over time. Generally speaking, there is typically a factor of 2×2$\times$2 × reduction in computation times with respect to the number of operations per second going from $64$-bit precision to $32$-bit precision, with considerably greater factors of reduction going to $16$-bit precision and to $8$-bit precision including as much as orders of magnitude factors of reduction; see, e.g., [23, 2, 1]. Then, since the cost of any additional iterations incurred by our approaches is often minimal in comparison with the significant reductions from mixed-precision computation in each iteration as shown in the above theoretical results and the empirical results in Section 6, our general classes of algorithmic approaches provide tremendous performance improvements over the most efficient existing methods.
@@
It is well understood that the computational performance of the advanced processor architectures of today can be hindered by memory bandwidth bottlenecks when the problem is sufficiently large relative to the available memory of the processor. In particular, rather than realizing the number of operations per second available from the processor in such cases, the computational performance is significantly reduced and instead dictated by the memory bandwidth of the processor architecture. Therefore, by exploiting reduced precision computation, our general classes of algorithmic approaches further enable us to handle significantly larger systems of linear equations in each iteration at the computational performance afforded by the processor architecture, including the mixed-precision reduction factors above. This is well beyond what is possible with existing methods whose performance would become memory-bandwidth limited for much smaller problem instances.
@@ 5 Proofs of Main Results
In this section, we provide the proofs of our main theoretical results presented in Section 4. The presentations of a couple of our proofs are reduced with complete proofs of such cases presented in Appendix A.
5.1 Proof of Lemma 4.1
@@
We exclude some of the steps and technical details of our proof due to space limitations, referring the reader to Appendix A.1 for such additional details.
From the specific supposition $\bm{\pi}^{(t-1)}_{i}=\bm{\pi}_{i}+O(\varphi)$ together with $10<z\leq\|\bm{\pi}_{i}\|_{1}<$ 1, we obtain
@@
Given the theorem statements $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}=\mathbf{s}\mathbf{R}$, we take the difference between the former statement and the latter statement on both sides and derive
@@
where (a) follows upon substituting (17) into (18) and multiplying both sides on the right by $\mathbf{S}$. Given the theorem statement $\|\mathbf{s}\|_{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ (according to the rightmost equation in (9)) and $\|\mathbf{S}\|=O(1)$ from the rightmost equation in (11), together with $\mathbf{R}^{(t-1)}=\mathbf{R}_{ij}+O(\varphi\epsilon)$ established above, we obtain
@@
It therefore follows from the last equation in (11) that $(\mathbf{s}^{(t-1)}-\mathbf{s})=O(\varphi)$, thus completing the proof. ∎
5.2 Proof of Theorem 4.2
Proof.
We exclude some of the steps and technical details of our proof due to space limitations, referring the reader to Appendix A.2 for such additional details.
From (1), we have $\bm{\pi}_{i}=\sum_{j=1}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}$ and $\bm{\pi}_{i}(\mathbf{I}-\mathbf{P}_{ii})=\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}$. Similarly, from (5) and (6), we obtain
@@
from which the desired result (16) follows.
Next, let the error in the approximate solution $\bm{\pi}^{(t-1)}$ at iteration $t-1$ of Algorithm 1 be of order $\varphi$, namely $\bm{\pi}^{(t-1)}=\bm{\pi}+O(\varphi)$. Then, from Lemma 4.1, we have
@@
It then follows from the initial arguments in the proof of Lemma 4.1 that
$\bm{\pi}^{(t)}=\bm{\pi}+O(\varphi\epsilon),\qquad\forall i\in[m],$
which yields the desired result. ∎
@@
Under the process employed for selecting the precision of computations in Step $\mathbf{5}$ and Step $\mathbf{3}$ of Algorithm 1 with respect to the condition number and the norm of the matrix $\mathbf{A}$ of the linear system, we know that the assumptions of the mixed-precision analysis of iterative refinement by Moler [20] are satisfied. It then follows from the convergence results of Moler [20] that, under these conditions, the iterative refinement method is guaranteed to converge linearly with an approximation error that decreases by a factor of $O(\kappa(\mathbf{A}))$ in each iteration, thus completing the proof. ∎
5.4 Proof of Lemma 4.4
Proof.
First, notice that the matrix $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ is block-lower triangular, with the $i$-th on-diagonal block being equal to $\mathbf{I}-\mathbf{P}_{ii}$. Since $\mathbf{P}_{ii}$ has only one eigenvalue that is close to one, and since this eigenvalue $\lambda_{i_{1}}$ is equal to $1-{g}_{i}\epsilon+o(\epsilon)$, the smallest eigenvalue of $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ is of the order $O(\epsilon)$. Expressing
@@
$\overline{\bm{\pi}}^{(t)\top}=\prod_{t=1}\left(\left(\left(\mathbf{D}^{\top}-\mathbf{U}^{\top}\right)^{-1}-\left((\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+1}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+2}+\cdots\right)\mathbf{D}^{-\top}\right)\mathbf{L}^{\top}\mathbf{C}\right)\bm{\pi}^{(0)\top}.$
Then, in order for the sequence $\{\overline{\bm{\pi}}^{(t)\top}\}$ to converge to the same limit as the sequence $\{\bm{\pi}^{(t)\top}\}$, we need to have $\prod_{t=1}(((\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}-((\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+1}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+2}+\cdots)\mathbf{D}^{-\top})\mathbf{L}^{\top}\mathbf{C})$ converge to the dominant eigendirection of $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}\mathbf{L}^{\top}\mathbf{C}$, which is true provided that $k_{t}$ increases monotonically with the number of outer-loop iterations $t$, thus completing the proof. ∎

### v2

@@ 4.1 Error and Convergence Analysis of Representative Conservative Example
Let $\mathbf{s}=\mathbf{s}\mathbf{R}$ and $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ with $\|\mathbf{s}\|_{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$, and suppose $\bm{\pi}^{(t-1)}$ is an $O(\varphi)$ approximation to $\bm{\pi}$, $t\in\mathbb{N}$, i.e., $\bm{\pi}^{(t-1)}_{i}=\bm{\pi}_{i}+O(\varphi)$. Then, we have
$\mathbf{s}^{(t-1)}-\mathbf{s}=O(\varphi),\qquad t\in\mathbb{N},$ (16)
@@
Further let $\mathbf{v}_{i}$ be the left eigenvector of $\mathbf{P}_{ii}$ corresponding to the exact invariant probability vector $\bm{\pi}_{i}$, $i\in[m]$. Then, there exists a constant ${b}_{i}=O(1)$ such that
$\|\bm{\pi}_{i}-{b}_{i}\mathbf{v}_{i}\|_{1}=O(\epsilon).$ (17)
We now establish our main theoretical results that show Algorithm 1 is guaranteed to converge and does so at a rate that decreases the approximation error of the estimate $\bm{\pi}^{(t)}$ of the exact solution $\bm{\pi}$ in each outer-loop iteration $t$ by a factor of $O(\epsilon)$. Based on our selection of the computational precision with respect to the condition number and the norm of the matrix of the linear system as defined in Section 3.2, these theoretical results also imply that our use of the IR method in the inner loop of Algorithm 1 is guaranteed to converge and does so at a rate that decreases the corresponding relative approximation error in each inner-loop iteration by a factor of $O(\kappa(\mathbf{A}))$.
@@
Consider Algorithm 1 employing the mixed-precision IR method under the precision-selection process described above in the inner loop Step $\mathbf{5}$ and Step $\mathbf{3}$ to compute the solution of the corresponding linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$. Then, supposing $\bm{\pi}^{(t-1)}$ is an $O(\varphi)$ approximation to $\bm{\pi}$, $t\in\mathbb{N}$, i.e., $\bm{\pi}^{(t-1)}=\bm{\pi}+O(\varphi)$, Algorithm 1 converges with an approximation error in the solution $\bm{\pi}^{(t)}$ at each outer-loop iteration $t$ that decreases by a factor of $O(\epsilon)$, i.e., $\bm{\pi}^{(t)}=\bm{\pi}+O(\varphi\epsilon)$ with $\epsilon<1$. Moreover, the IR method converges linearly with an approximation error that decreases by a factor of $O(\kappa(\mathbf{A}))$ in each inner-loop iteration.
The results of Lemma 17 and Theorem 4.2 are based on the assumption that the systems of linear equations in Step $\mathbf{5}$ and Step $\mathbf{3}$ are all solved at the exact accuracy of full precision. We note that the results in Theorem 4.2 on the outer-loop convergence rate associated with a factor of $O(\epsilon)$ and the inner-loop convergence rate associated with a factor of $O(\kappa(\mathbf{A}))$ hold in a similar manner for the alternative general classes of approximate mixed-precision computing methods of interest. We also note that the linear convergence rate factor of $O(\kappa(\mathbf{A}))$ for the inner-loop iterations is typically much faster than the convergence rate factor of $O(\epsilon)$ for the outer-loop iterations of Algorithm 1.
4.2 Error and Convergence Analysis of Representative Aggressive Example
Our main theoretical result is that Algorithm 2 — a representative aggressive example of our general algorithmic framework — is guaranteed to converge with an approximation error that decreases with respect to a tradeoff between an increase in the number of iterations $t$ of the outer loop for fewer numbers of RI steps $k_{t}$ in Step $\mathbf{5}$ and a decrease in the computational complexity of the RI method for smaller $k_{t}$. To consider these theoretical aspects of our second representative example, we begin by seeking to understand the effects of stopping the mixed-precision RI method after a certain number of steps, say $k\in\mathbb{N}$, as formally expressed in the following lemma.
@@ 4.2 Error and Convergence Analysis of Representative Aggressive Example
Consider computing the solution of the linear system $(\mathbf{D}^{\top}-\mathbf{U}^{\top})\mathbf{x}=\mathbf{b}$. Then, for any $k>0$, the absolute error satisfies
@@
Assuming that the dominant eigenvalue of the matrix $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}\mathbf{L}^{\top}\mathbf{C}$ is simple and has modulus $1$, the iterate $\bm{\pi}^{(t)}$ converges towards the corresponding eigenvector as formally expressed in the next lemma.
@@
Consider the update $\overline{\bm{\pi}}^{(t)\top}=((\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k_{t}]}^{-1}\mathbf{L}^{\top}\mathbf{C})\overline{\bm{\pi}}^{(t-1)\top}$ where $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}$ is applied via $k_{t}$ steps of the RI method, and $k_{t}$ is monotonically increasing with the number of outer-loop iterations $t$. Then, the sequence $\{\overline{\bm{\pi}}^{(t)}\}$ converges to the same limit as the sequence $\{\bm{\pi}^{(t)}\}$.
Our Algorithm 2 addresses a fundamental tradeoff between computational times on the one hand and error and convergence on the other hand. At one extreme, if $k_{t}$ grows too quickly with the outer-loop iterations $t\in\mathbb{N}$, then there will be error and convergence guarantees analogous to Theorem 4.2, but at the expense of little or no improvement in the computation times. At the other extreme, if $k_{t}$ grows too slowly with the outer-loop iterations $t\in\mathbb{N}$, then there will be significant improvements in the computation times, but at the expense of slow convergence and possibly even a lack of guaranteed convergence. The key to our Algorithm 2 is the enabling of a control sequence $k_{1},k_{2},\ldots$ that balances this fundamental tradeoff between the improvements in computation times and the speed of convergence.
The error $\mathbf{e}_{k}$ after $k$ steps of RI satisfies the equation $\mathbf{e}_{k}=(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top})^{k}\mathbf{e}_{0}$, where $\mathbf{e}_{0}$ denotes the initial approximation error. Let us now focus on the linear system $\mathbf{D}^{\top}-\mathbf{U}^{\top}\bm{\pi}^{(t)\top}=\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top}$. The RI method converges to the true solution $\bm{\pi}^{(t)\top}$ as long as the spectral radius $\rho(\cdot)$ of the matrix $\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top}$ satisfies $\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top})<1$ which is always the case when $\|\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top}\|<1$ since $\rho(\mathbf{A})\leq\|\mathbf{A}\|$ for any matrix $\mathbf{A}$. Therefore, after $k_{t}$ steps of RI, the norm of the approximation error $\bm{\pi}^{(t)\top}-\overline{\bm{\pi}}^{(t)\top}$ of Lemma 4.4 is reduced by a factor of $\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top})^{k_{t}}$ compared to the initial error. Hence, assuming that we know $\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top})$, we can reduce the initial error by a factor of $\hat{\epsilon}_{t}>0$ if we allow $k_{t}$ to be large enough. In particular, assuming that the initial approximation is always zero, this leads to
$k_{t}\geq\dfrac{\mathrm{ln}(\hat{\epsilon}_{t}/\|\bm{\pi}^{(t)\top}\|)}{\mathrm{ln}(\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top}))}.$
The above inequality tells us that, at iteration $t$ of Algorithm 2, we can determine the smallest number of iterations of RI in Step $\mathbf{5}$ that are required to achieve an arbitrary level of accuracy. For example, as used in our numerical experiments, we assume that $k_{1}$ is chosen as above and $k_{t}=2^{t}k_{0}$, i.e., the number of inner iterations increases geometrically. Starting with an inexact application of the RI method and increasing the number of RI steps $k_{t}$ with the number of outer-loop iterations $t$, Algorithm 2 is able to avoid over-solving in the earlier stages of Step $\mathbf{5}$ where the approximate stationary distribution $\bm{\pi}^{(t)}$ is far from $\bm{\pi}$. We note here that such tradeoffs are well known in numerical linear algebra and arise often when iterative solvers are used in shift-and-invert algorithms for eigenvalue computations [29].
@@
Consider Algorithm 2 employing the mixed-precision RI method under the $k_{t}$-selection process described above for the number of steps performed in the inner loop Step $\mathbf{5}$. Then, supposing $\bm{\pi}^{(t-1)}$ is an $O(\varphi)$ approximation to $\bm{\pi}$, $t\in\mathbb{N}$, i.e., $\bm{\pi}^{(t-1)}=\bm{\pi}+O(\varphi)$, Algorithm 2 converges with an approximation error in the solution $\bm{\pi}^{(t)}$ at each outer-loop iteration $t$ that decreases by a factor of $O(\epsilon)$, i.e., $\bm{\pi}^{(t)}=\bm{\pi}+O(\varphi\epsilon)$ with $\epsilon<1$, and the inner loop of Algorithm 2 at each outer-loop iteration $t$ converges to within an accuracy of $O(\epsilon^{t})$. Moreover, the RI method converges linearly with an approximation error that decreases by a factor of $O(\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top}))$ in each inner-loop iteration.
We note that the results in Theorem 4.5 on the outer-loop convergence rate associated with a factor of $O(\epsilon)$ and the inner-loop convergence rate associated with a factor similar to $O(\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top}))$ can hold in a related manner for appropriate alternatives in the general classes of approximate mixed-precision computing methods of interest. We also note that the linear convergence rate factor of $O(\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top}))$ for the inner-loop iterations is typically much faster, or can be made much faster by techniques such as preconditioning, than the convergence rate factor of $O(\epsilon)$ for the outer-loop iterations of Algorithm 2.
@@ 4.3 Mixed-Precision Computation Performance Analysis
Our general mathematical framework of numerical methods exploits a combination of advances in mixed-precision computer architectures and advances in iterative approximate computing approaches. Although well studied in machine learning and related applications, we note that such advances in mixed-precision computation have received far less consideration in the context of computing the stationary distribution of Markov processes. Beyond significantly reducing computation times, our use of lower-precision computation makes it possible to handle much larger systems of linear equations before hitting the memory bandwidth limitations of today’s highly-efficient advanced processors.
@@
More specifically, the performance tradeoff at the heart of our general algorithmic framework concerns, on the one hand, the significant reductions in execution times afforded by mixed-precision computation at the expense of inaccuracies in the results and, on the other hand, the ability to mitigate inaccurate computations, further reduce execution times and guarantee convergence afforded by iterative approximate computing methods. Sections 4.1 and 4.2 address the degree to which various representative examples of our algorithmic framework mitigate the impact of inaccurate computations, reduce execution times, and ensure convergence. The reductions in execution times provided by mixed-precision computation are also quite significant and vary from one computer processor architecture to the next, with increasing trends in execution-time reductions from mixed-precision computation over time. Generally speaking, there is typically a factor of 2×2$\times$2 × reduction in computation times with respect to the number of operations per second going from $64$-bit precision to $32$-bit precision, with considerably greater factors of reduction going to $16$-bit precision and to $8$-bit precision including as much as orders of magnitude factors of reduction; see, e.g., [30, 31, 32]. Then, since the cost of any additional iterations incurred by our representative examples is often minimal in comparison with the significant reductions from mixed-precision computation and iterative approximate approaches in each iteration as shown in the above theoretical results and the empirical results in Section 6, our general mathematical framework of numerical methods provides tremendous performance improvements over the most-efficient existing methods.
@@
It is well understood that the computational performance of the advanced processor architectures of today can be hindered by memory bandwidth bottlenecks when the problem is sufficiently large relative to the available memory of the processor. In particular, rather than realizing the number of operations per second available from the processor in such cases, the computational performance is significantly reduced and instead dictated by the memory bandwidth of the processor architecture. Therefore, by exploiting reduced-precision computation, our general algorithmic framework further enables us to handle significantly larger systems of linear equations in each iteration at the computational performance afforded by the processor architecture, including the mixed-precision reduction factors above. This is well beyond what is possible with existing methods whose performance would become memory-bandwidth limited for much smaller problem instances.
@@ 5 Proofs of Main Results
In this section, we provide the proofs of our main theoretical results presented in Section 4. The presentations of several of our proofs are shortened with complete proofs of such cases presented in Appendix A.
5.1 Proof of Lemma 17
@@
We exclude some of the steps and technical details of our proof to simplify the presentation, referring the reader to Appendix A.1 for such additional details.
From the specific supposition $\bm{\pi}^{(t-1)}_{i}=\bm{\pi}_{i}+O(\varphi)$ together with $10<\tau\leq\|\bm{\pi}_{i}\|_{1}<$ 1, we obtain
@@
Given the theorem hypotheses $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}=\mathbf{s}\mathbf{R}$, where the latter follows from (9), we take the difference between the former equation and the latter equation on both sides and derive
@@
where (a) follows upon substituting (18) into (19) and multiplying both sides on the right by $\mathbf{S}$. Given the theorem hypothesis $\|\mathbf{s}\|_{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ (according to the rightmost equation in (9)) and $\|\mathbf{S}\|=O(1)$ from the rightmost equation in (11), together with $\mathbf{R}^{(t-1)}=\mathbf{R}_{ij}+O(\varphi\epsilon)$ established above, we obtain
@@
It therefore follows from the last equation in (11) that $(\mathbf{s}^{(t-1)}-\mathbf{s})=O(\varphi)$, thus completing the proof of (16).
Lastly, from (1), we have $\bm{\pi}_{i}=\sum_{j=1}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}$ and $\bm{\pi}_{i}(\mathbf{I}-\mathbf{P}_{ii})=\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}$. Similarly, from (5) and (6), we obtain
@@
from which the desired result (17) follows. ∎
5.2 Proof of Theorem 4.2
Proof.
We exclude some of the steps and technical details of our proof to simplify the presentation, referring the reader to Appendix A.2 for such additional details.
Let the error in the approximate solution $\bm{\pi}^{(t-1)}$ at iteration $t-1$ of Algorithm 1 be of order $\varphi$, namely $\bm{\pi}^{(t-1)}=\bm{\pi}+O(\varphi)$. Then, from Lemma 17, we have
@@
It then follows from the initial arguments in the proof of Lemma 17 that
$\bm{\pi}^{(t)}_{i}=\bm{\pi}_{i}+O(\varphi\epsilon),\qquad\forall i\in[m],$
which yields the desired result.
Next, under the process employed for selecting the precision of computations in Step $\mathbf{5}$ and Step $\mathbf{3}$ of Algorithm 1 with respect to the condition number and the norm of the matrix $\mathbf{A}$ of the linear system, we know that the assumptions of the mixed-precision analysis of the IR method by Moler [26] are satisfied. It then follows from the convergence results of Moler [26] that, under these conditions, the IR method is guaranteed to converge linearly with an approximation error that decreases by a factor of $O(\kappa(\mathbf{A}))$ in each iteration, thus completing the proof. ∎
@@
First, notice that the matrix $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ is block-lower triangular, with the $i$-th diagonal block being equal to $\mathbf{I}-\mathbf{P}_{ii}$. Since $\mathbf{P}_{ii}$ has only one eigenvalue that is close to one, and since this eigenvalue $\lambda_{i_{1}}$ is equal to $1-{g}_{i}\epsilon+o(\epsilon)$, the smallest eigenvalue of $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ is of the order $O(\epsilon)$. Expressing
@@
$\overline{\bm{\pi}}^{(t)\top}=\prod_{t=1}(((\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}-((\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+1}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+2}+\cdots)\mathbf{D}^{-\top})\mathbf{L}^{\top}\mathbf{C})\bm{\pi}^{(0)\top}.$
Then, in order for the sequence $\{\overline{\bm{\pi}}^{(t)}\}$ to converge to the same limit as the sequence $\{\bm{\pi}^{(t)}\}$, we need to have
$\prod_{t=1}(((\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}-((\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+1}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+2}+\cdots)\mathbf{D}^{-\top})\mathbf{L}^{\top}\mathbf{C})$
converge to the dominant eigendirection of $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}\mathbf{L}^{\top}\mathbf{C}$, which is true provided that $k_{t}$ increases monotonically with the number of outer-loop iterations $t$, thus completing the proof. ∎